if two angles of a triangle are equal ant the measure of the thrd angle is 4 degress less than twice of eache of the other angels what is the measure of the largest angle?
Answer by rwm(914)

2x-4+2x=4x-4=180
2x-2=90
x-1=45
x=46
2x-4=92-4=88
46+46+88=180